# OpenStereo: A Comprehensive Benchmark for Stereo Matching and Strong Baseline

### 23 Dec 2023 v3

@@ Abstract
Stereo matching, a pivotal technique in computer vision, plays a crucial role in robotics, autonomous navigation, and augmented reality. Despite the development of numerous impressive methods in recent years, replicating their results and determining the most suitable architecture for practical application remains challenging. Addressing this gap, our paper introduces a comprehensive benchmark focusing on practical applicability rather than solely on performance enhancement. Specifically, we develop a flexible and efficient stereo matching codebase, called OpenStereo. OpenStereo includes training and inference codes of more than 12 network models, making it, to our knowledge, the most complete stereo matching toolbox available. Based on OpenStereo, we conducted experiments on the SceneFlow dataset and have achieved or surpassed the performance metrics reported in the original paper. Additionally, we conduct an in-depth revisitation of recent developments in stereo matching through ablative experiments. These investigations inspired the creation of StereoBase, a simple yet strong baseline model. Our extensive comparative analyses of StereoBase against numerous contemporary stereo matching methods on the SceneFlow dataset demonstrate its remarkably strong performance. The source code is available at https://github.com/XiandaGuo/OpenStereo.
@@ 1 Introduction
Stereo matching, a core subject in computer vision, focuses on calculating the disparity between the left and right images. It is vital in a wide array of applications such as robotics [1], autonomous driving [2, 3], and augmented reality [4], as it enables depth perception and 3D reconstruction of the observed scene.
@@
Traditional stereo matching algorithms typically rely on matching corresponding image regions between the left and right views based on their similarity measures. Several techniques have been proposed, including methods based on gray-level information [5, 6, 7], region-based approaches [8, 9], and energy optimization methods [10, 11]. Gray-level-based methods [5, 6, 7] compute disparities by matching pixels or groups of pixels with similar intensities in the left and right images. Region-based approaches [8, 9] group pixels into larger regions and then compute correspondences between these regions. Energy optimization methods [10, 11] involve minimizing an energy function that represents the disparity between image pairs to improve the accuracy of 3D reconstruction from stereo images. While these approaches have made noteworthy progress, they suffer from certain limitations and challenges, such as occlusions, textureless regions, and computational complexity.
@@
Recently, bolstered by extensive synthetic datasets [12, 13, 14, 15, 16], CNNs-based stereo matching methods [17, 18, 19, 20, 21, 22, 23, 24] has achieved impressive results. These methods have shown remarkable performance particularly in aspects of accuracy and efficiency. Unlike traditional methods, CNN-based stereo matching methods take left and right images as input and directly predict the disparity between them, which are end-to-end frameworks. In addition to the accuracy and efficiency benefits, CNN-based stereo matching methods have shown the potential for generalization to new environments and scenes beyond the training data. This is mainly due to the ability of CNNs to learn feature representations from raw data that are invariant to scale, rotation, and illumination changes. Furthermore, the emergence of large-scale datasets has paved the way for the integration of deep learning in stereo matching, allowing for more complex and sophisticated models to be trained.
Despite the considerable advancements made in CNN-based methods, the existing methods still face challenges in achieving both high accuracy and time efficiency. Encoder-decoder networks using 2D CNN methods [25, 12, 26, 27, 28, 24] offer efficiency but fall short in achieving extremely high precision, while the cost volume matching with 3D convolution methods [17, 18, 19, 29] deliver top-tier accuracy but pose challenges in deployment due to their substantial computational and memory requirements. However, choosing an appropriate architecture for stereo matching is still a challenging task, since the trade-off between accuracy and efficiency needs to be carefully balanced. In this context, recent studies have explored the potential of automated machine learning (AutoML) [39, 42, 23] on stereo matching, aiming to reduce the human effort required to design an efficient network structure. With the increasing number of new methods and datasets in stereo matching, there is a lack of a unified framework to support state-of-the-art models and datasets. To address these challenges, we propose OpenStereo, a platform for reproducing and comparing various stereo matching models on major datasets. By conducting extensive experiments, we aim to reproduce and assess the performance of various state-of-the-art models within a unified framework. Our platform also aims to facilitate the comparison of different models and provide a user-friendly environment for researchers to quickly reproduce and build new models.
@@
In summary, this paper presents substantial contributions to the domain of stereo matching research, focusing on three key aspects. (1) OpenStereo, a unified and extensible platform, enables researchers to conduct comprehensive stereo matching studies. (2) With the support of OpenStereo, we deeply revisit the recent stereo matching methods and significantly influence future research in this field. (3) We introduce StereoBase, a model notable for its structural simplicity and powerful performance, offering a fresh perspective for the design of upcoming stereo matching algorithms. These contributions collectively advance the state-of-the-art in stereo matching, offering both theoretical insights and practical tools to the research community.
@@ 2.1 Stereo Matching
With the rapid development of CNNs, there has been remarkable progress in the field of stereo matching. Based on the network pipeline of stereo matching, stereo matching methods can be roughly grouped into two categories [24], including the encoder-decoder network with 2D convolution (ED-Conv2D) and the cost volume matching with 3D convolution (CVM-Conv3D).
@@
Stereo Matching with CVM-Conv3D The CVM-Conv3D methods are proposed to improve the performance of depth estimation [17, 18, 20, 19, 30, 31, 23, 32, 33, 34, 3, 35, 36, 37, 29]. These techniques derive disparities from a 4D cost volume, primarily formed by concatenating left feature maps with their respective right analogs at each disparity level [18]. GCNet [17] firstly introduced a novel approach that combines 3D encoder-decoder architecture with a 2D convolutional network to obtain a dense feature representation, which is used to regularize a 4D concatenation volume. Following GCNet [17], PSMNet [18] proposes an approach for regularizing the concatenation volume by leveraging a stacked hourglass 3D convolutional neural network in tandem with intermediate supervision. To boost the expressiveness of the cost volume and thereby enhance performance in regions with ambiguity, GwcNet [19] proposes the group-wise correlation volume and ACVNet [37] proposes the attention concatenation volume. GANet [19] innovates with a semi-global aggregation layer and a local guided aggregation layer, designed to supplant the traditionally employed 3D convolutional layer. Based on GANet [19], DSMNet [30] constructs a robust “domain normalization” method to overcome the challenges in cross-domain generalization. CoEx [33] proposes a novel approach called Guided Cost volume Excitation (GCE), which leverages image guidance to construct a simple channel excitation of the cost volume. IGEV-Stereo [38] leverages an iterative geometry encoding volume to capture both local and non-local geometry information, which outperforms existing methods on KITTI benchmarks and achieves cross-dataset generalization and high inference efficiency. Besides, as a potential solution to reduce the need for humans to design a good volumetric method, Neural Architecture Search (NAS) [39, 40, 41] has been employed in stereo matching [23, 42].
However, these CVM-Conv3D methods continue to face challenges with time efficiency and substantial memory demands, hindering real-time inference, even on server-grade GPUs. Consequently, resolving these issues of accuracy and efficiency is crucial for their practical application in real-world scenarios.
@@
Stereo Matching with ED-Conv2D The ED-Conv2D methods [43, 12, 44, 45, 46, 47, 42, 24, 21, 48, 49, 50, 22, 3, 51], which adopt networks with 2D convolutions to predict disparity, has been driven by the need for improved accuracy, computational efficiency, and real-time performance. Subsequently, Mayer et al [12] present an end-to-end network, called DispNet, which is pure 2D CNN architectures. However, the model still faces challenges in capturing the matching features, resulting in poor estimation results. To overcome this challenge, the correlation layer is introduced in stereo matching [12, 52, 26, 27] to better capture the relationship between the stereo images. By incorporating this layer, the accuracy of the model is significantly improved. Furthermore, FADNet++ [24] proposes an innovative approach to efficient disparity refinement using residual learning [53] in a coarse-to-fine manner. AutoDispNet [42] applied neural architecture search to automatically design stereo matching network structures. AANet [21] leverages deformable convolution to enhance the standard convolution’s ability to model geometric transformations. More recently, Croco-Stereo [54] demonstrates that large-scale pre-training, when aligned with appropriately chosen pretext tasks, can significantly enhance stereo matching performance.
@@
These works represent the significant progress that has been made in stereo matching, highlighting the diverse range of methods and architectures that have been proposed to address the challenges associated with this problem.
@@ 2.2 Codebase
Numerous infrastructure code platforms have been developed in the deep learning research community, with the aim of facilitating research in specific fields. One such platform is OpenGait [55], a gait recognition library. OpenGait thoroughly examines the latest advancements in gait recognition, providing novel perspectives for subsequent research in this domain. In object detection, MMDetection [56] and Detectron2 [57] have emerged as an all-encompassing resource for several favored detection techniques. In Pose Estimation, OpenPose [58] has developed the first open-source system that operates in real-time for detecting the 2D pose of multiple individuals, including the detection of key points for the body, feet, hands, and face. With the rapid development of stereo matching, the prominence of an infrastructure code platform has risen markedly, underscoring the pressing demand for such a system.
@@ 3 OpenStereo
In recent times, the field of stereo matching has seen a surge in the development of new frameworks and evaluation datasets. Despite this growth, the absence of a unified and impartial evaluation platform in this domain stands out as a critical concern that warrants attention. To address this challenge and promote academic research and practical application we have developed OpenStereo, a PyTorch-based [59] toolbox that provides a reliable and standardized evaluation framework for stereo matching.
@@ 3.1 Design Principles of OpenStereo
As illustrated in Figure 2, OpenStereo covers the following standout features.
Compatibility with various frameworks. Currently, an increasing number of novel stereo matching methods are emerging in the field, such as Concatenatation-based [18], Correlation-based [21, 60, 24], Interlaced-based [3], Group-wise-correlation-based [38], Difference-based [44], and Combine-based methods [20]. As previously noted, many open-source methods are narrowly tailored to their specific models, making it challenging to extend to multiple frameworks. However, OpenStereo provides a solution to this problem by supporting all of the aforementioned frameworks. With OpenStereo, researchers and practitioners can easily compare and evaluate different stereo matching models under a standardized evaluation protocol.
Support for various evaluation datasets. OpenStereo serves as an all-encompassing tool, incorporating datasets frequently utilized by researchers in the field of stereo matching. OpenStereo not only supports synthetic stereo datasets such as SceneFlow [12], but also five real-world datasets: KITTI12 [15], KITTI15 [16], Middlebury [13], ETH3D [14], and DrivingStereo [61]. We introduce a suite of bespoke functions, meticulously crafted for each dataset, encompassing everything from initial data preprocessing to the final stages of evaluation.
Support for state-of-the-arts. OpenStereo successfully replicated various state-of-the-art stereo matching methods, including PSMNet [18], GwcNet [20], AANet [21], FADNet++ [24], CFNet [49], STTR [50], ACVNet [37], RAFT-Stereo [22], CoEx [33], CascadeStereo [31], MobileStereoNet [3] and IGEV [38]. Notably, the performance metrics we achieved, in most cases, surpass those reported in their original publications. This comprehensive replication effort provides a valuable resource for novices in the field, offering a rich array of official examples to facilitate an effective and efficient start in stereo matching research.
@@ 3.2 Main Modules
Technically, we align with the structure commonly found in PyTorch-based deep learning projects, organizing OpenStereo into three modules, data, modeling, and evaluation.
@@
Evaluation module is employed to meticulously assess and measure the performance of the developed model. Recognizing that different datasets require distinct evaluation protocols, we have incorporated these varied protocols directly into OpenStereo, significantly simplifying the evaluation process.
3.3 Training Pipeline
We propose a unified training pipeline with a hooking mechanism that can be customized to suit different training processes. Our pipeline is based on a common workflow that includes iterative training epochs and optional validation epochs. In each epoch, we forward and backward the model through multiple iterations. To increase flexibility and customization, we define a minimal pipeline that repeatedly forwards the model, and additional behaviors are implemented through a hooking mechanism. This mechanism allows users to define custom operations to be executed before or after specific steps in the training process. The pipeline defines several time points where users can register executable methods or hooks, such as before and after training epochs, before and after training iterations, and before and after validation epochs. These hooks are triggered at the specified time points according to their priority level, enabling users to customize their training processes in a more granular way. A typical training pipeline in OpenStereo is shown in Figure 3.
4 Benchmarks
OpenStereo contains high-quality implementations of popular stereo matching methods.
4.1 Supported Methods
PSMNet [18] is one of the most crucial and frequently cited networks in stereo matching. The official implementation code is available at https://github.com/JiaRenChang/PSMNet.
GwcNet [20] employs a Group-wise Correlation Stereo Network to exploit both local and global matching cost volumes. https://github.com/xy-guo/GwcNet.
AANet [21] focuses on adaptively fusing multi-scale features through a lightweight and efficient adaptive aggregation module. https://github.com/haofeixu/aanet.
FADNet++ [24] employs a two-stage architecture, consisting of a feature-metric aggregation stage and a disparity refinement stage. https://github.com/HKBU-HPML/FADNet.
CFNet [49] utilizes a collaborative feature learning strategy to enhance the representation capacity of features in the matching cost computation. https://github.com/gallenszl/CFNet.
STTR [50] focuses on geometric correspondence problems, including stereo matching, by employing transformers to capture long-range spatial and temporal context information. https://github.com/mli0603/stereo-transformer.
ACVNet [37] addresses the issue of constructing accurate and efficient cost volumes by adopting an asymmetric strategy. https://github.com/gangweiX/ACVNet.
RAFT-Stereo [22] builds upon the RAFT[62] architecture, originally designed for optical flow estimation. https://github.com/princeton-vl/RAFT-Stereo.
CoEx [33] focuses on context-aware features and cost volume aggregation. https://github.com/antabangun/coex.
CascadeStereo [31] adopts a cascaded architecture for accurate and efficient disparity estimation. https://github.com/alibaba/cascade-stereo.
MobileStereoNet [3] builds upon the StereoNet [44], which proposes Interlacing Cost Volume Construction. https://github.com/cogsys-tuebingen/mobilestereonet.
IGEV [38] builds a combined geometry encoding volume that not only encodes both geometric and contextual information but also intricately captures fine-grained details of local matching. https://github.com/gangweiX/IGEV.

### 08 Mar 2024 v4

@@ Abstract
Stereo matching aims to estimate the disparity between matching pixels in a stereo image pair, which is of great importance to robotics, autonomous driving, and other computer vision tasks. Despite the development of numerous impressive methods in recent years, replicating their results and determining the most suitable architecture for practical application remains challenging. Addressing this gap, our paper introduces a comprehensive benchmark focusing on practical applicability rather than solely on performance enhancement. Specifically, we develop a flexible and efficient stereo matching codebase, called OpenStereo. OpenStereo includes training and inference codes of more than 10 network models, making it, to our knowledge, the most complete stereo matching toolbox available. Based on OpenStereo, we conducted experiments and have achieved or surpassed the performance metrics reported in the original paper. Additionally, we carry out an exhaustive analysis and deconstruction of recent developments in stereo matching through comprehensive ablative experiments. These investigations inspired the creation of StereoBase, a strong baseline model. Our StereoBase ranks 1st on SceneFlow, KITTI 2015, 2012 (Reflective) among published methods and achieves the best performance across all metrics. In addition, StereoBase has strong cross-dataset generalization. Code is available at https://github.com/XiandaGuo/OpenStereo.
@@ 1 Introduction
Stereo matching is a fundamental topic in the field of computer vision, aiming to compute the disparity between a pair of rectified stereo images. It plays a crucial role in numerous applications such as robotics [60], autonomous driving [30, 39], and augmented reality [53], as it enables depth perception and 3D reconstruction of the observed scene.
@@
Traditional stereo matching algorithms typically rely on matching corresponding image regions between the left and right views based on their similarity measures. Several techniques have been proposed in the literature for stereo matching, including methods based on gray-level information [2, 23, 56], region-based approaches [62, 33], and energy optimization methods [37, 16]. Gray-level-based methods [2, 23, 56] compute disparities by matching pixels or groups of pixels with similar intensities in the left and right images. Region-based approaches [62, 33] group pixels into larger regions and then compute correspondences between these regions. Energy optimization methods [37, 16] formulate the stereo matching problem as an energy minimization problem, where the goal is to find the disparity map that minimizes a certain energy function. While these methods have achieved some success in stereo matching, they suffer from certain limitations and challenges, such as occlusions, textureless regions, and computational complexity.
@@
Recently, with the support of large synthetic datasets [28, 36, 38, 11, 29], CNNs-based stereo matching methods [19, 5, 58, 13, 52, 25, 7, 46] has achieved impressive results. These methods have shown remarkable performance in terms of accuracy and efficiency. Unlike traditional methods, CNN-based stereo matching methods are end-to-end frameworks that take a pair of stereo images as input and directly predict the disparity between them. In addition to the accuracy and efficiency benefits, CNN-based stereo matching methods have shown the potential for generalization to new environments and scenes beyond the training data. This is mainly due to the ability of CNNs to learn feature representations from raw data that are invariant to scale, rotation, and illumination changes. Furthermore, the availability of large-scale datasets has enabled the use of deep learning techniques for stereo matching, allowing for more complex and sophisticated models to be trained.
Nevertheless, we find that not all datasets are accompanied by official evaluation tools. For example, the DrivingStereo [55] dataset does not provide specific evaluation scripts, making comparative assessments challenging. The SceneFlow [28] dataset, with its finalpass and cleanpass data varieties, complicates fair model comparisons. Generalization experiments for stereo matching algorithms typically train on the SceneFlow dataset and evaluate on KITTI2012 [11], KITTI2015 [29], ETH3D [38], and Middlebury [36]. However, the inconsistency in evaluation data volume and metric definitions across studies leads to comparative discrepancies. Furthermore, even when source codes for models are available, reproducing the reported benchmarks is exceedingly challenging, e.g. STTR [24]. Hence, there’s a pressing need for a comprehensive benchmark study within the stereo matching community to enhance practicality and ensure consistent comparisons. To achieve this objective, we introduce a versatile stereo matching codebase: OpenStereo.
To promote scalability and adaptability, OpenStereo offers the following features: (1) Modual design, researchers can define a new model without the need to alter the model code itself by simply modifying a yaml configuration file. (2)Various frameworks, including Concatenation-based [5, 58], Correlation-based [52, 45, 46, 1], Interlaced-based [39], Group-wise-correlation-based [51], Combine-based methods [13], and Difference-based [21]. (3)Various datasets, including SceneFlow [28], KITTI2012 [11], KITTI2015 [29], Middlebury [36], ETH3D [38] and DrivingStereo [55] dataset. (4) State-of-the-art methods, including PSMNet [5], GwcNet [13], AANet [52], FADNet++ [46], CFNet [40], STTR [24], CoEx [1], CascadeStereo [12], MobileStereoNet [39] and IGEV [51].
Leveraging OpenStereo, we rigorously reassess various officially stated conclusions by meticulously re-implementing the ablation studies, including data augmentation, backbone architectures, cost construction, disparity regression and refinement processes. Based on the insights gleaned from these ablation experiments, we introduce StereoBase, a model that sets a new benchmark, surpassing recently proposed methods in terms of performance. Furthermore, we undertake a thorough investigation to establish that StereoBase is not only powerful but also serves as an empirically robust baseline model for stereo matching, demonstrating exceptional efficacy and resilience across diverse testing scenarios.
@@
In summary, this paper makes significant contributions to the field of stereo matching research from three key aspects. (1) OpenStereo, a unified and extensible platform, enables researchers to conduct comprehensive stereo matching studies. (2) With the support of OpenStereo, we conduct a profound revisitation and thorough deconstruction of recent stereo matching methodologies. (3) We introduce StereoBase, which sets a new benchmark with EPE of 0.34 on SceneFlow [28] and ranks 1st on KITTI2015 [29] and 2012(Reflective)[11] leaderboards among published methods. Furthermore, StereoBase exhibits exceptional cross-dataset generalization capabilities. These contributions collectively advance the state-of-the-art in stereo matching, offering both insights and practical tools to the stereo matching research community.
@@ 2.1 Stereo Matching
With the rapid development of CNNs, significant progress has been made in stereo matching. Based on the network pipeline of stereo matching, stereo matching methods can be roughly grouped into two categories [46], including the encoder-decoder network with 2D convolution (ED-Conv2D) and the cost volume matching with 3D convolution (CVM-Conv3D).
@@
Stereo Matching with CVM-Conv3D The CVM-Conv3D methods are proposed to improve the performance of depth estimation [19, 5, 13, 58, 59, 12, 7, 44, 1, 41, 39, 22, 26, 50, 31]. These methods learn disparities from a 4D cost volume, which is mainly constructed by concatenating left feature maps with their corresponding right counterparts across each disparity level [5]. GCNet [19] firstly introduced a novel approach that combines 3D encoder-decoder architecture with a 2D convolutional network to obtain a dense feature representation, which is used to regularize a 4D concatenation volume. Following GCNet [19], PSMNet [5] proposes an approach for regularizing the concatenation volume by leveraging a stacked hourglass 3D convolutional neural network in tandem with intermediate supervision. In order to enhance the expressiveness of the cost volume and ultimately improve performance in ambiguous regions, GwcNet [58] proposes the group-wise correlation volume and ACVNet [50] proposes the attention concatenation volume. GANet [58] proposes a semi-global aggregation layer and local guided aggregation layer to capture local and the whole-image cost dependencies respectively. Based on GANet [58], DSMNet [59] construct a robust “domain normalization” method to overcome the challenges in cross-domain generalization. CoEx [1] proposes a novel approach called Guided Cost volume Excitation (GCE), which leverages image guidance to construct a simple channel excitation of the cost volume. IGEV-Stereo [51] leverages an iterative geometry encoding volume to capture both local and non-local geometry information, which outperforms existing methods on KITTI benchmarks and achieves cross-dataset generalization and high inference efficiency. Besides, as a potential solution to reduce the need for humans to design a good volumetric method, Neural Architecture Search (NAS) [15, 27, 3] has been employed in stereo matching [7].
However, these CVM-Conv3D methods still suffer from low time efficiency and high memory requirements, which are far from real-time inference even on server GPUs. Therefore, it is essential to address the accuracy and efficiency problems for real-world applications.
@@
Stereo Matching with ED-Conv2D The ED-Conv2D methods [57, 28, 21, 34, 54, 9, 35, 46, 52, 43, 40, 24, 25, 39, 61], which adopt networks with 2D convolutions to predict disparity, has been driven by the need for improved accuracy, computational efficiency, and real-time performance. One of the early deep learning-based stereo matching methods, MC-CNN (Matching Cost CNN) [57], was proposed to learn a matching cost function for improving performance in the cost aggregation and optimization stages. Subsequently, Mayer et al. [28] present end-to-end networks for the estimation of disparity, called DispNet, which is pure 2D CNN architectures. However, the model still faces challenges in capturing the matching features, resulting in poor estimation results. To overcome this challenge, the correlation layer is introduced in the end-to-end architecture [28, 8, 17, 18] to better capture the relationship between the left and right images. By incorporating this layer, the accuracy of the model is significantly improved. Furthermore, FADNet++ [46] proposes an innovative approach to efficient disparity refinement using residual learning [14] in a coarse-to-fine manner. AutoDispNet [35] applied neural architecture search to automatically design stereo matching network structures. More recently, Croco-Stereo [47] shows that large-scale pre-training can be successful for stereo matching through well-adapted pretext tasks. This method can achieve state-of-the-art performance without using task-specific designs, like correlation volume or iterative estimation.
@@
These works represent the significant progress that has been made in the field of stereo matching, highlighting the diverse range of methods and architectures that have been proposed to address the challenges associated with this problem.
@@ 2.2 Codebase
Numerous infrastructure code platforms have been developed in the deep learning research community, with the aim of facilitating research in specific fields. One such platform is OpenGait [10], a gait recognition library. OpenGait thoroughly examines the latest advancements in gait recognition, providing novel perspectives for subsequent research in this domain. In object detection, MMDetection [6] and Detectron2 [49] have emerged as an all-encompassing resource for several favored detection techniques. In pose estimation, OpenPose [4] has developed the first open-source system that operates in real-time for detecting the 2D pose of multiple individuals, including the detection of key points for the body, feet, hands, and face. In stereo matching, it is noteworthy that not all datasets are accompanied by official evaluation tools. For instance, the DrivingStereo [55] dataset does not have official evaluation codes, and there is a lack of unified tools for assessing the model generalization across different domains. This absence of standardized evaluation resources contributes to the observed discrepancies in cross-domain evaluations of the same model as reported in different studies. Therefore, it is time to build a benchmark for stereo matching.
@@ 3 OpenStereo
In recent years, there has been a proliferation of new frameworks and evaluation datasets for stereo matching. However, the lack of a unified and fair evaluation platform in this field is a significant issue that cannot be ignored. To address this challenge and promote academic research and practical application we have developed OpenStereo, a pyTorch-based [32] toolbox that provides a reliable and standardized evaluation framework for stereo matching.
@@ 3.1 Design Principles of OpenStereo
As shown in Figure 3, our developed OpenStereo covers the following highlight features.
Modular Design. OpenStereo adopts a modular design, greatly facilitating researchers in exploring new networks. By simply modifying a YAML configuration file, researchers can define a new model without the need to alter the model code itself. This design significantly lowers the barriers for researchers to extend or integrate additional algorithms and modules within the framework. This approach empowers researchers to freely combine and customize their algorithms with minimal code composition, enhancing the framework’s usability and adaptability.
Compatibility with various frameworks. Currently, more and more stereo matching methods have emerged, such as Concatenation-based [19, 5, 58, 7], Correlation-based [52, 45, 46, 1], Group-wise-correlation-based [51], Difference-based [21], Interlaced-based [39], and Combine-based methods [13, 40]. As mentioned above, many open-source methods have a narrow focus on their specific models, making it challenging to extend to multiple frameworks. However, OpenStereo provides a solution to this problem by supporting all of the aforementioned frameworks. With OpenStereo, researchers and practitioners can easily compare and evaluate different stereo matching models under a standardized evaluation protocol.
Support for various evaluation datasets. OpenStereo is a comprehensive tool that includes datasets commonly used by researchers in stereo matching. OpenStereo not only supports synthetic stereo datasets such as SceneFlow [28], but also five real-world datasets: KITTI2012 [11], KITTI2015 [29], ETH3D [38], Middlebury [36], and DrivingStereo [55]. We introduce a suite of bespoke functions, meticulously crafted for each dataset, encompassing everything from initial data preprocessing to the final stages of evaluation. The evaluation module includes the submission of the results to all benchmarks.
Support for state-of-the-arts. In our work, we have successfully replicated various state-of-the-art stereo matching methods, including PSMNet [5], GwcNet [13], AANet [52], FADNet++ [46], CFNet [40], STTR [24], CoEx [1], CascadeStereo [12], MobileStereoNet [39] and IGEV [51]. Notably, the performance metrics we achieved, in most cases, surpass those reported in their original publications. This comprehensive replication effort provides a valuable resource for novices in the field, offering a rich array of official examples to facilitate an effective and efficient start in stereo matching research.
@@ 3.2 Main Modules
Technically, we follow the design of most PyTorch deep learning projects and divide OpenStereo into three modules, data, modeling, and evaluation.
@@
Evaluation module is utilized to assess the performance of the obtained model. Recognizing that different datasets require distinct evaluation protocols, we have incorporated these varied protocols directly into OpenStereo, significantly simplifying the evaluation process.
4 Revisit Deep Stereo Matching
4.1 Recheck on Previous Methods
PSMNet [5] is one of the most crucial and frequently cited networks in stereo matching. The official implementation code is available at https://github.com/pytorch/vision. GwcNet [13] employs a Group-wise Correlation Stereo Network to exploit both local and global matching cost volumes (https://github.com/xy-guo/GwcNet). AANet [52] focuses on adaptively fusing multi-scale features through a lightweight and efficient adaptive aggregation module ( https://github.com/haofeixu/aanet). FADNet++ [46] employs a two-stage architecture, consisting of a feature-metric aggregation stage and a disparity refinement stage. (https://github.com/HKBU-HPML/FADNet). CFNet [40] utilizes a collaborative feature learning strategy to enhance the representation capacity of features in the matching cost computation (https://github.com/gallenszl/CFNet). STTR [24] focuses on geometric correspondence problems, including stereo matching, by employing transformers to capture long-range spatial and temporal context information (https://github.com/mli0603/stereo-transformer). CoEx [1] focuses on context-aware features and cost volume aggregation (https://github.com/antabangun/coex). CascadeStereo [12] adopts a cascaded architecture for accurate and efficient disparity estimation, whose official code is available at (https://github.com/alibaba/cascade-stereo). MobileStereoNet [39] builds upon the StereoNet [21], which proposes Interlacing Cost Volume Construction (https://github.com/cogsys-tuebingen/mobilestereonet). IGEV [51] builds a combined geometry encoding volume that not only encodes both geometric and contextual information but also intricately captures fine-grained details of local matching (https://github.com/gangweiX/IGEV).